Problem 9-11 Valuation of a constant growth stock A stock is expected to pay a dividend of $1.00 the end of the year (that is, D1 = $1.00), and it should continue to grow at a constant rate of 4% a year. If its required return is 1496, what is the stocks expected price 3 years from today? Round your answer to two decimal places.
Answer #1

Answer:

Problem 9 - 11

What is the stock's expected price 3 years from today?

Current Price = D1 / (ke - g) = 1/(0.14-0.04) = $10

Price after 3 years = 10(1 + 0.04)^3 = $11.25

Microtech Corporation 1.5 D 4 1.25 (1.36) 1.70 D5-1.70 (1.36)2.312 Compute D 6-2315 (1.05)-2.4276 just to use for constant growth model (DVM) for valuing dividends from year 6 to infinity P 5-D 6 / (kg)-2.4276 / (0.18-0.05)-$80.92